Everything is working pretty well. This morning the pressure was about 1.2e-6 Torr. I connected the high voltage amplifier and I could drive the disk without problems.
I measured the beam shape and size at the QPD. We have about 50 uW, we see a TEM01-like mode due to the interference of the two disk surfaces (this is normal). The beam is about 3 mm in diameter. using this information and the estimated optical lever length of 1.2 m, I calbrated the QPD NORM signals in units of angular motion of the disk surface. The computation posted in CRIME_Lab/60 is actually wrong. I'll post the correct one later.
Injecting broadband white noise I could excite all the modes that are visible up to about 30kHz. I tuned the COMSOL model, by changing the thickess of the disk to 1.017 mm, to fit the frequency of the first few modes. Here are the modes I could measure:
Take a look at the attached PDF file for the shape of all the modes, including all that are not visible. We see all the modes we expect to be able to excite with the central suspension of the disk.
The roughing pump is making a lot of non stationary low frequency noise. I turned it off, and the pressure stayed constant at 1.2e-6 Torr over about 1.5 hours. Here's the difference in the QPD spectrum:
It turned out that I have enough excitation authority to knock the disk out of the right place. So I had to vent to recover the situation. I'll open the chamber tomorrow and see what happened.
Here's a first bird eye look at the ring downs. We see beating of the two almost denegerate modes in some cases. Fits will follow, using the procedure I used for the LMA measurements.